OpenStudy (anonymous):

x=y-1 x+2y=8 solve by elimination method

OpenStudy (anonymous):

.x-y=-1 x+2y=8 subtract ---------- 0-3y=-9 y=9/3=3 ...ans sub this to either of the two eq say eq1 x-y=-1 x-3=-1 y=3-1=2 ans

OpenStudy (anonymous):

Thanks